OpenStudy (jellybot23):

How would you turn this equation for an ellipse 2(x+4)^2 + 3(y-1)^2 = 24 into standard form??

OpenStudy (tkhunny):

Which Standard form is that? Perhaps multiply out the squares, collect like terms, and you will be close?

OpenStudy (jellybot23):

Well isn't standard form (x-h)^2 / a^2 + (y-k)^2 / b^2 = 1?

OpenStudy (jellybot23):

@tkhunny

OpenStudy (tkhunny):

Yes, that is a standard form for an ellipse. That's why I asked. Just divide by 24 to get that standard form, There is also the standard quadratic form: \(ax^{2} + bxy + cy^{2} + dx + ey + f = 0\) In our case, b = 0, but everything else might be there.

OpenStudy (jellybot23):

Would I have to simplify because there is a 2 in front of (x+4) and a 3 in front of (y-1)? @tkhunny

OpenStudy (tkhunny):

Just divide by 24. \(\dfrac{2x^{2}}{24} = \dfrac{x^{2}}{12}\)

OpenStudy (jellybot23):

Okay that's what I thought, but I wanted to make sure! I haven't dealt with ellipse equations yet this year where the bottom number wasn't a Perfect square, so I was a little disoriented haha Thank you!

OpenStudy (tkhunny):

Next time, SHOW what you thought. It makes for more pleasant conversation. :-)